# Exploring Remote Collaboration: The Impact of Avatar Representation on Dyadic Haptic Interactions in Shared Virtual Environments

### v1

@@ Abstract
This study is the first to explore the interplay between haptic interaction and avatar representation in Shared Virtual Environments (SVEs). We focus on their combined effect on social presence and task-related scores in dyadic collaborations. In a series of experiments, participants performed the plate control task with haptic interaction under four avatar representation conditions: avatars of both participant and partner were displayed, only the participant’s avatar was displayed, only the partner’s avatar was displayed, and no avatars were displayed. The study finds that avatar representation, especially of the partner, significantly enhances the perception of social presence, which haptic interaction alone does not fully achieve. In contrast, neither the presence nor the type of avatar representation impacts the task performance or participants’ force effort of the task, suggesting that haptic interaction provides sufficient interaction cues for the execution of the task. These results underscore the significance of integrating both visual and haptic modalities to optimize remote collaboration experiences in virtual environments, ensuring effective communication and a strong sense of social presence.
@@
Avatar Representation, Haptic Interaction, Remote Collaboration
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@
Virtual Environment (VE) technologies have provided unprecedented new experiences to users. In particular, Virtual Reality (VR) devices, represented by recent advances in head-mounted displays (HMDs), have become more performant and affordable, making VR increasingly accessible to the general public. In VEs, users can represent body movements through avatar embodiment[1]. Various avatar representations have been reported to alter user experiences[2]. Accurate measurement of self-body and reflection in an avatar can enhance the sense of ownership of the body[3]. Furthermore, accurate body tracking enhances the user experience in VEs[4].
@@
Shared virtual environments (SVEs), where VE is shared with other users, allow users to share space despite each user being remotely located. SVEs are attracting attention as a new communication tool.[5]. In SVEs, visual interaction, which is crucial for users to communicate with each other[6], enhances social presence by providing the feeling of “being there”[7].
@@
Several studies have reported that the absence of an avatar representation can reduce social presence[8, 9], a decrease in social presence can potentially lead to reduced social communication, which is a critical issue. An avatar is one of the critical elements that provide a visual modality in non-verbal communication. Moreover, remote collaboration using SVEs is expected[10]. Remote collaboration through SVEs allows you to solve tasks collaboratively without location constraints, potentially solving issues such as labor shortages.
@@
In recent years, VR applications that integrate haptic technologies have attracted attention. Haptic feedback can provide information that cannot be conveyed visually and is an essential modality for humans[11]. The haptic modality is crucial when touching objects or moving, but it also builds a beneficial communication channel during interactions with others[12, 13]. In SVEs, haptic interaction provides various benefits to users. This means that haptic interactions influence not only physical performance but also the psychological user experience. Some studies have found that haptic interactions contribute to users feeling a higher social presence[14, 15, 16, 17, 18]. Therefore, haptic interaction is indispensable for performing physical tasks in remote collaboration.
Currently, it is still unclear whether displaying avatars is necessary for physical remote collaboration. To our knowledge, no study has investigated the impact of avatar representation on users when there is haptic interaction. Does haptic interaction alone provide sufficient social presence, or is the representation of avatars also necessary? We found no clues to this question.
@@
Therefore, this paper compares the representation of avatars in dyadic physical remote collaboration under four conditions (both the user’s and the partner’s avatars are displayed, only the user’s avatar is displayed, only the partner’s avatar is displayed, and no avatars are displayed). Social presence, task performance, and users’ force effort were measured. The results showed that participants felt a decrease in social presence when the partner’s avatar was not displayed. Meanwhile, task performance and effort remained constant under all conditions. Therefore, participants found the social presence provided by haptic interactions to be insufficient, but the task execution was not affected by the representation of avatars.
@@
Based on these results, we derive the necessity of avatar representation for developing physical remote collaboration. The contributions of this paper are as follows: 1) When the purpose is solely task execution in physical remote collaboration, the representation of avatars can be omitted. 2) In social communication channels, visual interaction through avatar representation is more important than haptic interaction.
I RELATED WORK
I-A Avatar representation in shared virtual environment
@@
In social interactions, visual information plays a crucial role. Especially in SVEs provided through HMDs, users expect a bodily experience that mimics real-world movements and synchronizes with their own actions. Avatars, as a form of visual information, not only offer users a realistic representation of their bodies, but also facilitate the presence of others[19]. The presence of avatars of others, compared to audio and video communication channels, sufficiently enhances the user’s experience of emotional intimacy[6]. The representation of avatars plays a crucial role in determining user experience and task performance[20]. Gao et al. reported that the use of full-body avatars improved task performance compared to using only hands or hands and torso[21]. Moreover, Pan et al. conducted competitive and cooperative tasks in virtual environments with and without self-avatars and in face-to-face settings[10]. They found that the absence of a self-avatar during cooperative tasks led to a decrease in task scores. However, no significant differences were observed in competitive tasks. Thus, the impact of avatar representation on performance may vary depending on the context of the task.
@@
I-B Social presence and haptic interaction in shared virtual environment
@@
Haptic interaction allows users to share forces among themselves, enabling them to perform physical remote collaboration. Physically connected dyads form communication channels through the exchange of forces, allowing them to receive information from their partner [12]. Haptic interactions offer various physical benefits to dyads: improvement in time to motor learning [24, 25], enhancement of performance [26], maintenance of balance [27], and negotiation of strategies [12].
@@
Haptic interactions also provide user experience benefits beyond physical advantages. Several studies have investigated the impact of haptic interactions in SVEs by toggling the presence and absence of such interactions among dyads [14, 15, 16, 17, 18]. Chellali et al. conducted a task in which dyads trained in injection procedures using a Follow-Follower system and demonstrated that haptic interactions improved performance and co-presence[15]. Additionally, Sallnas et al. found that exchanging objects through haptic interactions enhanced social presence[14]. In summary, haptic interactions provide both physical and psychological benefits to users.
Both visual interaction through avatar representation and haptic interaction have been confirmed to enhance social presence and task performance. The question arises: Can social presence and task performance be maintained through only haptic interactions when avatar representation is lost? To date, no studies have focused on both avatar representation and haptic interactions simultaneously. Our study is the first to address this question.
II EXPERIMENTAL SETUP
@@
We developed an experimental environment for a physical remote collaboration task as shown in Fig. 1. The SVE displayed a plate and avatars representing the participants’ body movements. The dyes manipulated the plate through a haptic device, performing cooperative tasks.
@@
Figure 1: Experimental setup. (a) The shared virtual environment. (b) Participants could perform the cooperative task through the independent haptic interface. (c) Participant install the HMD and see the SVE. (d) The plate control task.
II-A Shared virtual environments
@@
Our developed SVE was created using Unity 2022.3.23f1. In this SVE, a plate was fixed in a floating position at a height of 1 m, surrounded by avatars. Dyad’s avatars were positioned face-to-face across the plate. The plate is a square with each side 0.4 m. Participants accessed the SVE using the Meta Quest 3, which tracked the position of participants’ heads, wrists, and each joint of the fingers, applying these positions in real time to the avatars’ actions. Data from Meta Quest 3 were captured using the Meta XR Core SDK v63.0.0. The movements of the avatar’s body parts not derived from captured positions were determined using Final IK. Inverse Kinematics (IK) techniques can cause errors, which can cause a feeling of discomfort for participants[28]. However, since the participants moved within a limited range in this study, these errors were not significant enough to affect the experiment and were not be considered.
The avatars used were generic with no gender or distinctive features, reflecting only the participants’ height as physical information. This choice was made to eliminate any impact on the facial expressions or gender of the participants. In fact, the force exerted on an avatar has been reported to decrease when it appears to be female[29]. Furthermore, participants interacting with expressive avatars feel stronger social presence and attractiveness and show superior task performance compared to those interacting with non-expressive avatars[30]. We needed to remove these variables to study primitive effects.
II-B Haptic Interface
@@
Haptic interactions between dyads were bilateral. One of these haptic interfaces is provided for each participant and operates independently. The haptic interface featured two motors and had a gimbal mechanism that could rotate along the roll and pitch axes. Thus, participants were capable of two degrees of freedom of movement. The motors used were Mitsubishi Electric’s HG-MR43, and to maintain high backdrivability, they were driven directly. Direct drive is often adopted in the research of physical Human–Human Interaction (pHHI) [31, 13]. Extending from the gimbal mechanism was a rod that could be gripped with both hands. The length of the rod from the left to the right was 0.4 m, and the length to the gimbal mechanism was 0.2 m, perfectly corresponding to the length of the plate. The gimbal mechanism was fixed at a height of 1 m and the participants operated it through the haptic interface in a standing position.
II-C Haptic interaction between dyad
@@
To provide haptic interaction to the dyad, the study applied acceleration-based admittance control to the haptic interfaces, allowing bilateral driving [32]. This form of control, suitable for building haptic interactions, allows precise transmission of forces among participants [33]. A disturbance observer was used to implement acceleration-based control [34], which suppresses external forces in the control system, aligning the command acceleration with the controlled acceleration. Furthermore, a reaction torque observer was employed to measure participants’ forces without a force sensor [35].
@@
Admittance control defines how a system responds to human operator forces through its dynamics, allowing designers to arbitrarily design these dynamics. This capability allows haptic interactions to be provided as if the dyads are jointly manipulating a single object in reality, a method employed in pHHI studies [12]. In this research, the dynamics of the plate were shared among the dyads.
@@
Here, the dyad’s force inputs through the haptic interfaces produced the forces that actuated the plate:
@@
where $\tau_{1}$ and $\tau_{2}$ represent each participant’s force input, with p and r indicating the pitch and roll axes, respectively.
The dynamics of the plate are expressed as:
@@
where $\ddot{\bm{\theta}}^{plt}$ and $\dot{\bm{\theta}}^{plt}$ represent the plate’s angular acceleration and velocity vectors, respectively, integrated to obtain the plate’s angular vector $\bm{\theta}^{plt}$. $\ddot{\bm{\theta}}^{plt}$, $\dot{\bm{\theta}}^{plt}$ and $\bm{\theta}^{plt}$ are
@@
$\theta$ and $\phi$ stand for pitch and roll axes angle, respectively. These dynamics, with $M$ and $D$ representing the plate’s mass and viscous coefficient, respectively, used values of 3.28 and 1.5 in the experiment.
II-D Plate control task

### v2

# Exploring Remote Collaborative Tasks: The Impact of Avatar Representation on Dyadic Haptic Interactions in Shared Virtual Environments

@@ Abstract
This study is the first to explore the interplay between haptic interaction and avatar representation in Shared Virtual Environments (SVEs). Specifically, how these factors shape users’ sense of social presence during dyadic collaborations, while assessing potential effects on task performance. In a series of experiments, participants performed the collaborative task with haptic interaction under four avatar representation conditions: avatars of both participant and partner were displayed, only the participant’s avatar was displayed, only the partner’s avatar was displayed, and no avatars were displayed. The study finds that avatar representation, especially of the partner, significantly enhances the perception of social presence, which haptic interaction alone does not fully achieve. However, neither the presence nor the type of avatar representation impacts the task performance or participants’ force effort of the task, suggesting that haptic interaction provides sufficient interaction cues for the execution of the task. These results underscore the significance of integrating both visual and haptic modalities to optimize remote collaboration experiences in virtual environments, ensuring effective communication and a strong sense of social presence.
@@
HCI, Avatar Representation, Haptic Interaction, Haptic communication, Remote Collaboration, Virtual human, Social interaction, Collaborative virtual environments, Shared virtual environments
I Introduction
@@
Virtual Environment (VE) technologies provide unprecedented experiences to users. In particular, Virtual Reality (VR) devices, represented by recent advances in head-mounted displays (HMDs), have become performant and affordable, making VR increasingly accessible to the general public. In VEs, users can represent their body movements through their avatars[1]. Various avatar representations have been reported to alter user experiences[2]. Accurate measurement of self-body and reflection in an avatar can enhance the sense of ownership of the body[3]. Furthermore, accurate body tracking enhances the user experience in VEs[4].
@@
Shared virtual environments (SVEs), where VE is shared with other users, allow users to share space despite each user being remotely located. SVEs are attracting attention as a new communication tool [5]. In SVEs, visual interaction, which is crucial for users to communicate with each other[6], enhances social presence by providing the feeling of “being there”[7].
@@
Several studies have reported that the absence of an avatar representation can reduce social presence[8, 9], a decrease in social presence can potentially cause reduced social communication, which is a critical issue. An avatar is a critical element that provide a visual modality in non-verbal communication. Moreover, remote collaboration using SVEs is expected[10]. Remote collaboration through SVEs allows you to solve tasks collaboratively without location constraints, potentially solving issues, such as labor shortages.
@@
Recently, VR applications that integrate haptic technologies have attracted attention. Haptic feedback can provide information that cannot be conveyed visually and is an essential modality for humans[11]. The haptic modality is crucial when touching objects or moving; however, it builds a beneficial communication channel during interactions with others[12, 13]. In SVEs, haptic interaction provides various benefits to users. Therefore, haptic interactions influence physical performance as well as the psychological user experience. Some studies have found that haptic interactions contribute to users feeling a higher social presence[14, 15, 16, 17, 18]. Therefore, haptic interaction is indispensable for performing physical tasks in remote collaboration.
In the course of our research aimed at enabling physical collaborative tasks through remote means, we developed an SVE. This work enabled us to determine how essential the effects of avatar representation might be in physical remote collaboration. Although numerous studies have reported that displaying avatars in SVEs can enhance both user experience and physical performance, we wondered if this impact persists when haptic interaction is present. In particular, in physical remote collaboration, one’s presence is constantly conveyed to others through the haptic modality, raising the question of whether avatar representation remains influential under such circumstances. Investigating the influence of avatars is important for designing remote collaborative tasks, as it provides valuable guidance for developers and designers.
Nevertheless, it remains unclear whether the display of avatars is essential for physical remote collaboration. To our knowledge, no study has systematically examined the effect of avatar representation on user experience in the context of haptic interaction. Specifically, we inquire: Can haptic interaction alone generate sufficient social presence, or is visual avatar representation also required? We found no empirical evidence that directly addresses this question.
@@
Therefore, in this study, we systematically compare avatar representation in dyadic physical remote collaboration under four conditions (i.e. both the user’s and the partner’s avatars are displayed, only the user’s avatar is displayed, only the partner’s avatar is displayed, and no avatars are displayed). We measured social presence, task performance, and users’ force effort. The results showed that participants felt a decrease in social presence when the partner’s avatar was not displayed. Meanwhile, task performance and effort remained constant under all conditions. These findings suggest that although haptic interaction alone does not fully compensate for the absence of visual cues in terms of social presence, it does not hinder task completion. Consequently, participants found the social presence provided by haptic interactions to be insufficient; however, the task execution was not affected by the representation of avatars.
In this study, we used a generic, gender-neutral avatar to remove confounding variables, such as voice or distinctive features, thereby isolating the effect of avatar presence on the interplay between visual and haptic cues. However, this minimalistic approach may limit the generalizability of our findings to scenarios involving realistic or personalized avatars with additional communication cues (e.g., voice or facial expressions).
@@
Based on these results, we emphasize the necessity of avatar representation for developing physical remote collaboration especially for maintaining social presence. The contributions of this paper are as follows: 1) If the purpose is solely task execution in physical remote collaboration, avatar representation can be omitted without compromising performance. 2) In social communication contexts, visual interaction through avatar representation is more critical than haptic interaction for maintaining social presence.
II RELATED WORK
II-A Avatar representation in shared virtual environment
@@
In social interactions, visual information plays a crucial role. Especially in SVEs provided through HMDs, users expect a bodily experience that mimics real-world movements and synchronizes with their own actions. Avatars, as a form of visual information, offer users a realistic representation of their bodies, as well as facilitate the presence of others[19]. The presence of avatars of others, compared to audio and video communication channels, sufficiently enhances the user’s experience of emotional intimacy[6]. The representation of avatars plays a crucial role in determining user experience and task performance[20]. Gao et al. reported that the use of full-body avatars improved task performance compared to using only hands or hands and torso[21]. Moreover, Pan et al. conducted competitive and collaborative tasks in VEs with and without self-avatars and in face-to-face settings[10]. They found that the absence of a self-avatar during collaborative tasks caused a decrease in task scores. However, no significant differences were observed in competitive tasks. Thus, the impact of avatar representation on performance may vary depending on the context of the task.
@@
II-B Social presence and haptic interaction in shared virtual environment
@@
Haptic interaction allows users to share forces among themselves, enabling them to perform physical remote collaboration. Physically connected dyads form communication channels through the exchange of forces, allowing them to receive information from their partner [12]. Such interactions offer various physical benefits to dyads: improvement in time to motor learning [24, 25], enhancement of performance [26], maintenance of balance [27], and negotiation of strategies [12].
@@
They also allow user to experience benefits beyond physical advantages. Several studies have investigated the impact of haptic interactions in SVEs by toggling the presence and absence of such interactions among dyads [14, 15, 16, 17, 18]. Chellali et al. conducted a task in which dyads trained in needle insertion procedures using a Follow–Follower system, and demonstrated that haptic interactions improved performance and co-presence[15]. Additionally, Sallnas et al. found that exchanging objects through haptic interactions enhanced social presence[14].
Therefore, haptic interactions provide both physical and psychological benefits to users, including improved motor learning, better performance, and enhanced social presence. In particular, the social presence fostered by haptic interactions supports the idea that they can serve as a communication channel in social interaction contexts. Meanwhile, previous findings indicate that visual interaction through avatar representation enhances social presence and task performance.
Focusing on physical remote collaboration using an SVE, haptic communication is indispensable and realized through haptic feedback, whereas avatar representation likely serves as the visual communication channel. Additionally, verbal communication may be involved. Thus, in SVE-based physical remote collaboration, social interaction is supported concurrently by multiple channels. Consequently, a critical question arises: do performance and social presence depend on a single channel, all channels, or specific combinations of channels? Furthermore, if one channel is lost, do performance and social presence decline, or are they preserved through the others? To the best of our knowledge, no research has specifically examined the impact of avatar representation in the context of physical remote collaboration. Our study is the first to address this question.
III METHODS
III-A Research hypothesis
This study investigates the impact of avatar representation in remote physical collaboration. In particular, previous studies have highlighted the roles of haptic interaction and task context in shaping user experience, social presence, and performance in SVEs or remote settings. Therefore, we propose the following hypotheses to elucidate how different forms of avatar representation might affect various aspects of collaborative work—even when participants are physically apart.
  • H1: Previous studies in SVE have reported that social presence can be heightened by perceiving the existence of others through haptic modalities [14, 15, 16, 17, 18]. This may be equivalent to the social presence arising from linguistic communication or visual interaction (e.g., avatars) [8] that have been noted in prior work. In physical remote collaboration using SVE, even if an avatar is not displayed, users can continue to feel their partner’s presence through haptic interaction. Therefore, regardless of the avatar representation, users can maintain the same level of social presence.
  • H2: It has been reported that the effect of avatar representation on users differs depending on the task context in which they engage within SVE [10]. In a context that requires intensive focus on task execution, the user is primarily devoted to achieving performance, thereby diminishing the influence of avatar representation. In physical remote collaboration, where the goal is to accomplish a particular task, the dyad may continue to achieve their highest possible performance regardless of avatar representation. Consequently, performance remains constant irrespective of the avatar representation.
  • H3: In physical collaborative tasks, it has been reported that force is used to perform the task as well as to negotiate intentions with one’s partner [12]. Considering this social role of force, the psychological effects induced by changes in avatar representation may ultimately alter the way force is applied.
III-B Experimental design
  • ALL: Both participant’s and partner’s avatars were displayed in the VE. It assumed that both users had body tracking devices.
  • OTHER: Only the partner’s avatar was displayed.
The conditions ALL, ALONE, and NONE are those anticipated in real environments. However, OTHER has been included based on our previous research [28], but it is a condition that does not occur in reality (as only the partner receives body tracking information, and no avatar is displayed on one’s own screen). We were interested in the effects that occur when self-avatar is not displayed. If self-avatar alone is not displayed in the SVE, can participants maintain their sense of social presence? In addition, each condition was applied simultaneously to both participants.
III-C Apparatus and Implementation
@@
We developed an experimental environment for a physical remote collaboration task as shown in Fig. 2. The SVE displayed a plate and avatars representing the participants’ body movements. The dyads manipulated the plate through a haptic device, performing collaborative tasks.
@@
Figure 2: Experimental setup. (a) The shared virtual environment. (b) Participants could perform the collaborative task through the independent haptic interface. (c) Participant install the HMD and see the SVE. (d) The plate control task.
III-C1 Shared virtual environments
@@
Our developed SVE was created using Unity 2022.3.23f1. In this SVE, a plate was fixed in a floating position at a height of 1 m, surrounded by avatars. Dyad’s avatars were positioned face-to-face across the plate. The plate is a square with each side 0.4 m. Participants accessed the SVE using the Meta Quest 3, which tracked the position of participants’ heads, wrists, and each joint of the fingers, applying these positions in real time to the avatars’ actions. Data from Meta Quest 3 were captured using the Meta XR Core SDK v63.0.0. The movements of the avatar’s body parts not derived from captured positions were determined using Final IK. Inverse Kinematics (IK) techniques can cause errors, ultimately causing a feeling of discomfort for participants[29]. However, because the participants moved within a limited range in this study, these errors were not sufficiently significant to affect the experiment; thus, they were not be considered.
The employed avatars were generic with no gender or distinctive features, reflecting only participants’ height as physical information. Their overall skin color was uniform, exhibiting a cream-like hue that falls between yellow and white, and textures were smooth with no variation across body parts. In addition, there were no prominent features in the chest or groin area, making it difficult to visually identify a particular gender. Thus, the influence of sex or facial personalization was minimized. However, the force exerted on an avatar has been reported to decrease when it appears female [30]. Furthermore, participants interacting with expressive avatars perceive stronger social presence and attractiveness, exhibiting better task performance than those with non-expressive avatars [31]. We adopted this minimalistic design to isolate fundamental effects of avatar representation and haptic interaction, avoiding confounds introduced by realistic or expressive avatars.
III-C2 Haptic Interface
@@
Haptic interactions between dyads were bilateral. One of these haptic interfaces is provided for each participant and operates independently. The haptic interface featured two motors, and had a gimbal mechanism that could rotate along the roll and pitch axes. Thus, participants could manage two degrees of freedom of movement. The motors used were Mitsubishi Electric’s HG-MR43, and to maintain high backdrivability, they were driven directly. Direct drive is often adopted in the research of physical Human–Human Interaction (pHHI) [32, 13]. Extending from the gimbal mechanism was a handle that could be gripped with both hands. The length of the handle from the left to the right was 0.4 m while that to the gimbal mechanism was 0.2 m, perfectly corresponding to the length of the plate. The gimbal mechanism was fixed at a height of 1 m and the participants operated it through the haptic interface in a standing position.
III-C3 Haptic interaction between dyad
@@
To provide haptic interaction to the dyad, we applied acceleration-based admittance control to the haptic interfaces, allowing bilateral driving [33]. This form of control, suitable for establishing haptic interactions, enables precise force transmission between participants [34]. Bilateral control represents motion transfer by ensuring that actuators satisfy Newton’s third law of action and reaction, as well as positional consistency [35]. Owing to these motion transfer properties, bilateral control has been widely adopted in various teleoperation systems [36, 37, 38]. To achieve both force and positional transmission with a single actuator, force-based and position-based motions must be integrated at the acceleration level [39]. In this control scheme, the desired acceleration is provided to the control system as the command, and the system’s response acceleration must match this target acceleration. We employed a disturbance observer to implement acceleration-based control [40], which suppresses external forces in the control system, aligning the command acceleration with the controlled acceleration. In addition, a reaction torque observer was used to measure participant forces without a force sensor [41].
@@
Admittance control defines how a system responds to human operator forces through its dynamics, thereby allowing designers to freely specify these dynamics. This capability allows for it to provide haptic interactions as though the dyads are jointly manipulating a single object in the real world, a method used in pHHI studies [12]. In this study, the dynamics of the plate were shared between the two participants.
@@
Here, the forces that each dyad member applied through their respective haptic interfaces generated the actual force acting on the plate:
@@
where $\tau_{1}$ and $\tau_{2}$ denote each participant’s force input, with the superscripts $p$ and $r$ indicating the pitch and roll axes, respectively.
The plate’s dynamics are expressed as:
@@
where $\ddot{\bm{\theta}}^{plt}$ and $\dot{\bm{\theta}}^{plt}$ represent the plate’s angular acceleration and velocity vectors, respectively, which are integrated to obtain the plate’s angular position vector $\bm{\theta}^{plt}$. These vectors are defined as follows:
@@
where $\theta$ and $\phi$ are the angles around the pitch and roll axes, respectively. In the experiment, the mass $M$ and the viscous coefficient $D$ of the plate were set to 3.28 and 1.5, respectively.
Through admittance control, participants can jointly manipulate a single shared object, determining its motion by the forces they apply—just as objects behave in the real world. If no haptic feedback is provided (i.e., no haptic interaction), the plate’s position reflects the average of each participant’s manual inputs, causing a discrepancy between the visually perceived plate position and participants’ actual movements. Such conditions are not appropriate for executing physical collaborative tasks. Therefore, in this study, we consistently provided haptic feedback to ensure realistic interaction.
III-C4 Plate control task